An object weighs 294.9N on a planet where the acceleration due to gravity is 4.9m/s². What is its weight (in N) on earth?

1
Expert's answer
2022-01-10T13:48:19-0500

Let's first find the mass of the object:


"m=\dfrac{W}{g}=\dfrac{294.9\ N}{4.9\ \dfrac{m}{s^2}}=60.18\ kg."

Finally, we can find the weight of the object on Earth:


"W=mg=60.18\ kg\times9.8\ \dfrac{m}{s^2}=589.76\ N."